# Context is Key: A Benchmark for Forecasting with Essential Textual Information

### v1

@@ E.1 Scaling for cross-task aggregation
In an ideal scenario, all instances of a tasks would be fully independent. In that case then Eq. 2 would not contain any information about the target time series in the benchmark instances, making the RCPRS a proper scoring rule. However, due to possible overlaps in the time windows used when creating the instances and to auto-correlations, we cannot guarantee independence between instances, and thus we cannot guarantee that the RCPRS is actually a proper scoring rule. Note that this deviation from a proper scoring rule is minor, and has a much smaller effect than the one due to the common approach of normalizing the CRPS using the Mean Absolute Value of the ground-truth.
@@ E.2 CRPS and twCRPS
We get the threshold-weighted CRPS (twCRPS) from Eq. 4 by adding a weighting function $w(x)$ to it:
@@
Eq. 7 can readily be generalized to a multivariate forecast, by using any $\mathbb{R}^{d}\rightarrow\mathbb{R}$ chaining function.
@@ E.3 Estimating the CRPS using samples
Computing the CRPS using Eq. 3 or Eq. 4 directly would be extremely hard for most of the baselines included in our experiments. Instead, it is more computationally convenient to use an estimator of the CRPS which uses a finite number of samples $x_{1}$, …, $x_{M}$ from the forecasting distribution. An unbiased estimator of the CRPS created from Eq. 4 is:
@@
A faster estimator which gives the same result as Eq. 8 (up to numerical accuracy) is the one based on the probability weighted moment form of the CRPS (Taillardat et al., 2016; Zamo & Naveau, 2018):
@@
where the $x_{n}$ have been sorted in ascending order. We used Eq. 9 in our metric, since it is as accurate as Eq. 8, while only having a $O(M\log M)$ time complexity.
@@ E.5 Covariance of two CRPS estimators
where the sums are over the various samples $n$ and $x_{i}$ and $x_{j}$ and the ground-truth values.
@@
$\mathrm{Cov}\left(\text{CRPS}\!\!\left(\widetilde{X}_{i},x_{i}\right),\text{CRPS}\!\!\left(\widetilde{X}_{j},x_{j}\right)\right)=\\ \begin{aligned} -\frac{1}{M}&\operatorname*{\mathbf{E}}_{\widetilde{X}_{i}}|\widetilde{X}_{i}-x_{i}|\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime}_{j}}|\widetilde{X}^{\prime}_{j}-x_{j}|\\ +\frac{1}{M}&\operatorname*{\mathbf{E}}_{\widetilde{X}_{i}}|\widetilde{X}_{i}-x_{i}|\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime}_{j}}\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime\prime}_{j}}|\widetilde{X}^{\prime}_{j}-\widetilde{X}^{\prime\prime}_{j}|\\ +\frac{1}{M}&\operatorname*{\mathbf{E}}_{\widetilde{X}_{i}}\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime}_{i}}|\widetilde{X}_{i}-\widetilde{X}^{\prime}_{i}|\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime\prime}_{j}}|\widetilde{X}^{\prime\prime}_{j}-x_{j}|\\ -\frac{2M-3}{2M(M-1)}&\operatorname*{\mathbf{E}}_{\widetilde{X}_{i}}\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime}_{i}}|\widetilde{X}_{i}-\widetilde{X}^{\prime}_{i}|\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime\prime}_{j}}\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime\prime\prime}_{j}}|\widetilde{X}^{\prime\prime}_{j}-\widetilde{X}^{\prime\prime\prime}_{j}|\\ +\frac{1}{M}&\operatorname*{\mathbf{E}}_{(\widetilde{X}_{i},\widetilde{X}_{j})}|\widetilde{X}_{i}-x_{i}|\cdot|\widetilde{X}_{j}-x_{j}|\\ -\frac{1}{M}&\operatorname*{\mathbf{E}}_{(\widetilde{X}_{i},\widetilde{X}_{j})}\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime}_{i}}|\widetilde{X}_{i}-\widetilde{X}^{\prime}_{i}|\cdot|\widetilde{X}_{j}-x_{j}|\\ -\frac{1}{M}&\operatorname*{\mathbf{E}}_{(\widetilde{X}_{i},\widetilde{X}_{j})}\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime}_{j}}|\widetilde{X}_{i}-x_{i}|\cdot|\widetilde{X}_{j}-\widetilde{X}^{\prime}_{j}|\\ +\frac{1}{2M(M-1)}&\operatorname*{\mathbf{E}}_{(\widetilde{X}_{i},\widetilde{X}_{j})}\operatorname*{\mathbf{E}}_{(\widetilde{X}^{\prime}_{i},\widetilde{X}^{\prime}_{j})}|\widetilde{X}_{i}-\widetilde{X}^{\prime}_{i}|\cdot|\widetilde{X}_{j}-\widetilde{X}^{\prime}_{j}|\\ +\frac{M-1}{M(M-1)}&\operatorname*{\mathbf{E}}_{(\widetilde{X}_{i},\widetilde{X}_{j})}\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime}_{i}}\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime\prime}_{j}}|\widetilde{X}_{i}-\widetilde{X}^{\prime}_{i}|\cdot|\widetilde{X}_{j}-\widetilde{X}^{\prime\prime}_{j}|,\\ \end{aligned}$
@@
To get an estimate of covariance using our $M$ samples, we can estimate each of these terms using their respective unbiased estimators. Once we have compute an estimate of the variance for a single task instance, the overall variance for a full task is computed using the formula for the variance of the average of multiple independent variables. One slight disadvantage of using this method, is that it offers now guarantee that the RCPRS variance estimate will be non-negative, so in the rare cases where the estimate for the variance of a full task is negative, we clip it to 0.

### v2

@@ E.1 Scaling for cross-task aggregation
In an ideal scenario, all instances of a tasks would be fully independent. In that case then Equation 2 would not contain any information about the target time series in the benchmark instances, making the RCPRS a proper scoring rule. However, due to possible overlaps in the time windows used when creating the instances and to auto-correlations, we cannot guarantee independence between instances, and thus we cannot guarantee that the RCPRS is actually a proper scoring rule. Note that this deviation from a proper scoring rule is minor, and has a much smaller effect than the one due to the common approach of normalizing the CRPS using the Mean Absolute Value of the ground-truth.
@@ E.2 CRPS and twCRPS
We get the threshold-weighted CRPS (twCRPS) from Equation 4 by adding a weighting function $w(x)$ to it:
@@
Equation 7 can readily be generalized to a multivariate forecast, by using any $\mathbb{R}^{d}\rightarrow\mathbb{R}$ chaining function.
@@ E.3 Estimating the CRPS using samples
Computing the CRPS using Equation 3 or Equation 4 directly would be extremely hard for most of the baselines included in our experiments. Instead, it is more computationally convenient to use an estimator of the CRPS which uses a finite number of samples $x_{1}$, …, $x_{M}$ from the forecasting distribution. An unbiased estimator of the CRPS created from Equation 4 is:
@@
A faster estimator which gives the same result as Equation 8 (up to numerical accuracy) is the one based on the probability weighted moment form of the CRPS (Taillardat et al., 2016; Zamo & Naveau, 2018):
@@
where the $x_{n}$ have been sorted in ascending order. We used Equation 9 in our metric, since it is as accurate as Equation 8, while only having a $O(M\log M)$ time complexity.
@@ E.5 Covariance of two CRPS estimators
where the sums are over the various samples $n$ and $x_{i}$ and $x_{j}$ are the ground-truth values.
@@
$\mathrm{Cov}\left(\text{CRPS}\!\!\left(\widetilde{X}_{i},x_{i}\right),\text{CRPS}\!\!\left(\widetilde{X}_{j},x_{j}\right)\right)=$ $-\frac{1}{M}\operatorname*{\mathbf{E}}_{\widetilde{X}_{i}}|\widetilde{X}_{i}-x_{i}|\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime}_{j}}|\widetilde{X}^{\prime}_{j}-x_{j}|$
$+\frac{1}{M}\operatorname*{\mathbf{E}}_{\widetilde{X}_{i}}|\widetilde{X}_{i}-x_{i}|\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime}_{j}}\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime\prime}_{j}}|\widetilde{X}^{\prime}_{j}-\widetilde{X}^{\prime\prime}_{j}|$
$+\frac{1}{M}\operatorname*{\mathbf{E}}_{\widetilde{X}_{i}}\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime}_{i}}|\widetilde{X}_{i}-\widetilde{X}^{\prime}_{i}|\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime\prime}_{j}}|\widetilde{X}^{\prime\prime}_{j}-x_{j}|$
$-\frac{2M-3}{2M(M-1)}\operatorname*{\mathbf{E}}_{\widetilde{X}_{i}}\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime}_{i}}|\widetilde{X}_{i}-\widetilde{X}^{\prime}_{i}|\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime\prime}_{j}}\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime\prime\prime}_{j}}|\widetilde{X}^{\prime\prime}_{j}-\widetilde{X}^{\prime\prime\prime}_{j}|$
$+\frac{1}{M}\operatorname*{\mathbf{E}}_{(\widetilde{X}_{i},\widetilde{X}_{j})}|\widetilde{X}_{i}-x_{i}|\cdot|\widetilde{X}_{j}-x_{j}|$
$-\frac{1}{M}\operatorname*{\mathbf{E}}_{(\widetilde{X}_{i},\widetilde{X}_{j})}\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime}_{i}}|\widetilde{X}_{i}-\widetilde{X}^{\prime}_{i}|\cdot|\widetilde{X}_{j}-x_{j}|$
$-\frac{1}{M}\operatorname*{\mathbf{E}}_{(\widetilde{X}_{i},\widetilde{X}_{j})}\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime}_{j}}|\widetilde{X}_{i}-x_{i}|\cdot|\widetilde{X}_{j}-\widetilde{X}^{\prime}_{j}|$
$+\frac{1}{2M(M-1)}\operatorname*{\mathbf{E}}_{(\widetilde{X}_{i},\widetilde{X}_{j})}\operatorname*{\mathbf{E}}_{(\widetilde{X}^{\prime}_{i},\widetilde{X}^{\prime}_{j})}|\widetilde{X}_{i}-\widetilde{X}^{\prime}_{i}|\cdot|\widetilde{X}_{j}-\widetilde{X}^{\prime}_{j}|$
$+\frac{M-1}{M(M-1)}\operatorname*{\mathbf{E}}_{(\widetilde{X}_{i},\widetilde{X}_{j})}\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime}_{i}}\operatorname*{\mathbf{E}}_{\widetilde{X}^{\prime\prime}_{j}}|\widetilde{X}_{i}-\widetilde{X}^{\prime}_{i}|\cdot|\widetilde{X}_{j}-\widetilde{X}^{\prime\prime}_{j}|,$
@@
To get an estimate of the covariance using our $M$ samples, we can estimate each of these terms using their respective unbiased estimators. Once we have computed an estimate of the variance for a single task instance, the overall variance for a full task is computed using the formula for the variance of the average of multiple independent variables. One slight disadvantage of using this method, is that it offers no guarantee that the RCPRS variance estimate will be non-negative, so in the rare cases where the estimate for the variance of a full task is negative, we clip it to 0.